OpenStudy (anonymous):

Can anyone help me with this: p-32=31/4h-5? I don't understand.

OpenStudy (anonymous):

what does the question ask? does it say solve? simplify?

OpenStudy (anonymous):

It says simplify

OpenStudy (anonymous):

oh, then it is easy, you just got to get the like terms together

OpenStudy (anonymous):

or, you can solve for either p or h if you solve for p you will get: p = (31/4)*h+27 and if you solve for h you will get: h = -108/31+(4/31)*p

OpenStudy (anonymous):

Fractions are so hard for me to understand.

OpenStudy (anonymous):

I see how you got p=31/4+27 but what do you do from there?